OpenStudy (anonymous):

On a sunny day, sunrays which propagate in the water make an angle of 50◦ with the water surface. How high (in degrees) is the Sun above the horizon?

OpenStudy (anonymous):

Use Snell's Law. The refractive index between one medium and another is:\[ \frac{\sin I}{\sin R}\] The refractive index of water to air is about 1.33